OpenStudy (anonymous):

Two right triangles are graphed on a coordinate plane. One triangle has a vertical side of 4 and a horizontal side of 10. The other triangle has a vertical side of 12 and a horizontal side of 30. Could the hypotenuses of these two triangles lie along the same line? No, because they are not similar triangles No, because one is larger than the other Yes, because they are similar triangles Yes, because they are both right triangles

OpenStudy (ineedhelpfast12):

C because 4 and 10 were multiplied by 3, making 12 and 30, which makes the triangle 3 times bigger

OpenStudy (anonymous):

yeah that makes sense thanks :)

OpenStudy (ineedhelpfast12):

actually, it is D

OpenStudy (ineedhelpfast12):

same reasoning though

OpenStudy (anonymous):

i put that and my teacher said it was wrong i think

OpenStudy (ineedhelpfast12):

If you don't mind, what grade are you in?

OpenStudy (ineedhelpfast12):

then yes it is indeed C. I hate second guessing myself

OpenStudy (anonymous):

Sorry about that my wifi was out but I'm in 8th :)

OpenStudy (anonymous):

Can you help with another?

OpenStudy (ineedhelpfast12):

me too, my teacher decided to moves us onto Algebra 1, and sure!

OpenStudy (anonymous):

A line passes through (2, 8) and (4, 12). Which equation best represents the line? y = 1 over 2 x + 6 y = 2x + 7 y = 2x + 4 y = 1 over 2 x + 10

OpenStudy (anonymous):

I can tell, you're smart lol.

OpenStudy (ineedhelpfast12):

thank you!

OpenStudy (ineedhelpfast12):

y=2x+4

OpenStudy (anonymous):

Which equation best represents the line? y = 1 over 3. x − 1 y = 3x − 1 y = −x + 1 over 3. y = 3x + 1 http://prntscr.com/6zpncb

OpenStudy (ineedhelpfast12):

y=3x-1

OpenStudy (anonymous):

can you explain how you got those

OpenStudy (ineedhelpfast12):

okay. so the line starts at (0,-1), so -1 is the y-intercept. to find the slope, you just go up from (0,-1) to (1,2), so the line went up from 0 to 1, which goes on the bottom, and from -1 to 2, so 3 would be on the top. since there is a 1 on the bottom of the 3, you can just simplify to be 3x

OpenStudy (anonymous):

I got it!

OpenStudy (ineedhelpfast12):

yay!

OpenStudy (anonymous):

lol thanks again!

OpenStudy (ineedhelpfast12):

the swag master strikes again!

OpenStudy (anonymous):

that's weird.. but I'll fan you on my other account :P

OpenStudy (ineedhelpfast12):

okay (think of me saying it like that guy from the okay vine

OpenStudy (anonymous):

I'm not on vine very much so I can't relate xD

OpenStudy (ineedhelpfast12):

aww pooh. :( oh well, nice chatting with you. I hope to hear from you soon, so I can help thee

OpenStudy (ineedhelpfast12):

sorry to intrude, but does your username have anything to do with dirty Dan from spongebob?

OpenStudy (anonymous):

Yes.